# A Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor

### v2

@@ Abstract
The strict geometric rules that define aperiodic tilings lead to the unique spectral and transport properties of quasicrystals, but also limit our ability to design them. In this work, we explore the first continuously tunable family of two-dimensional aperiodic tilings in which the underlying real-space geometry becomes a control knob of the wave-function’s quantum geometric tensor. The real-space geometry can be used to tune into topological phases occupying an expanded phase space compared to crystals, or into a disorder-driven topological Anderson insulator. The quantum metric can also be tuned continuously, opening new routes towards tunable single- and many-body physics in aperiodic solid-state and synthetic systems.
Introduction—The electronic properties of quasicrystals can be strikingly different to those of crystals [1, 2, 3]. For example, quasicrystals exhibit rotational symmetries forbidden in crystals, and their wave-functions are neither localized nor extended [1, 2, 3]. However, the strict mathematical rules that determine the quasicrystalline atomic arrangements and the special wave-functions they host also restricts the possibility of tuning them. Different realizations of two-dimensional quasicrystals, such as the Penrose [4] or the Ammann-Beenker [5] tilings, cannot be smoothly deformed into one another. Hence, continuously tuning the electronic properties of quasicrystals is typically limited to changing model parameters, as in crystals [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
Figure 1: Tiles and tilings for different values of the tile side length, defined by two side lengths, $\ell$ and $1-\ell$. (a) shows a selection of tiles, including the Chevron, Hat, Turtle, and Comet. Their corresponding aperiodic tilings are shown in (b-e). For each $\ell$ we place two orbitals $l$ at every vertex, with onsite terms $\epsilon_{ll^{\prime}}$. Pairs of sites $\{\alpha,\beta\}$ are connected by the space and orbital dependent hopping $T_{ll^{\prime}}(\mathbf{r}_{\alpha\beta})$, see Eq. (A Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor).
@@
Here we show that a newly discovered family of aperiodic tilings [66, 67] offers a real-space geometric pathway to tune the quantum geometric tensor. This is the only family of aperiodic tilings parametrically connected by a single real parameter $\ell$  [^1] Aperiodic tilings can be generated by the cut and project method. By varying continuously the cut angle one might expect to realize a continuous family of aperiodic tilings. However, varying this angle continuously typically results in jumps when new tiles appear, see e.g. this video. We thank F. Flicker for this remark.. This parameter determines the position of all atomic sites while preserving the quasicrystalline long-range order. Allowing the hoppings to depend on atomic positions, we show that changes in $\ell$ translate into changes of the quantum geometric tensor [69, 29, 32], expressed in terms of a real-space Chern marker [70] and quantum metric [71, 72]. Changes in the real-space Chern marker induced by $\ell$ induce phase transitions in and out of a topological phase, which exists in an expanded parameter space compared to the same model in crystals. The parameter $\ell$ also tunes the trace of the quantum metric [26, 27, 28, 29, 30], which contributes to the superfluid stiffness and kinetic inductance [73, 74, 75, 25, 76, 77], enters optical sum rules [78, 79, 80], and determines localization [81] and entanglement properties [82]. Hence, changing the real-space geometry using $\ell$ allows one to tune the quantum geometric tensor.
@@
Figure 2: Spectral and topological properties of the model for different parameter values $\ell$. (a) Energy spectrum of the bulk Hamiltonian as a function of the parameter $\ell$. We obtain the bulk spectrum by excluding states whose weight inside the yellow contour in (d) and (e) is smaller than their weight outside it. Magenta dashed lines indicate the $\ell$ values used in panels (b) and (c). The yellow dashed line indicates the Fermi energy $E_{F}$. (b) Spectral function $\mathcal{A}(E,{\bf p})$ for $\ell=0.23$, showing a full gap at the energy indicated by the white dashed line. (c) Spectral function $\mathcal{A}(E,{\bf p})$ for $\ell=1/3$, showing the presence of edge states within bulk gaps. (d) Local Chern marker $\mathcal{C}_{{\bf r}_{i}}$ Eq. (2), evaluated at the energy indicated by the white dashed line in (c) for $\ell=1/3$. The yellow solid line outlines the bulk region used to perform the averaging. In this region, the average value of the local Chern marker is $0.98$. (e) Local quantum metric $g_{\bf r}$, Eq. (5), computed at the $E_{F}$ indicated by the dashed line in (c) for $\ell=1/3$. We observe that the local quantum metric takes large values at the edges, indicating delocalized edge states. In all the panels the system has $2530$ sites and $M/t=-2.7$. In panels (b) and (c) we used the kernel polynomial method [83] to compute the spectral function with 512 moments.
@@
Geometric control of topology—Fig. 1a shows the Hat family of aperiodic tilings, parameterized by a real number $\ell\in\left[0,1\right]$. It parametrizes the two independent lengths of parallel sides, $a=\ell$ and $b=1-\ell$. By tuning $\ell$, one can continuously interpolate between different tiles in the family. The Hat tiling [$\ell=1/(1+\sqrt{3})$] made headlines as the first aperiodic monotile [66, 67]. Other named tiles are the Specter [$\ell=1/2$], the Turtle [$\ell=\sqrt{3}/(1+\sqrt{3})$], the Chevron [$\ell=0$], and the Comet [$\ell=1$] which sets our unit of length. Previous work has shown that a non-bipartite version of the Hat tiling shares similar electron properties with those of graphene [84] when all the hoppings are set to be equal. Here, we focus on the bipartite version of these tilings, and describe their construction in Appendix A of the Supplemental Material (SM) [85], by adapting the code published along with Ref. [86].
@@
Figure 3: Topological phase diagram. (a) Bulk half-signature of the spectral localizer at $E_{0}=E_{F}$ as a function of both the geometrical parameter $\ell$ and the value of the onsite term $M/t$. $E_{F}$ is fixed at $3/4$ filling out of a total of $N_{\mathrm{sites}}\times N_{\mathrm{orbs}}=2530\times 2$ states. We observe topological phases for a wider parameter region compared to the square-lattice crystalline model, whose phase boundaries are delimited by the black dashed lines. (b) Localizer gap corresponding to (a). The gap in the spectrum of the spectral localizer closes at the boundary between the topological and trivial phases of panel (a). (c) Bulk half-signature of the spectral localizer averaged over 50 disorder realizations at $E_{0}=E_{F}$ and $M/t=0.5$, as a function of $\ell$ and onsite disorder strength $W/t$. We observe two types of behavior as we increase $W/t$ for fixed $\ell$: either the clean system is topological and transitions into a trivial Anderson insulator (e.g. at $\ell=0.5$), or a trivial clean system transitions into a topological Anderson insulator, to later become a trivial Anderson insulator (e.g. at $\ell=0.8$ or $\ell=0.2$). (d) The localizer gap vanishes at the interface between the topological and trivial Anderson insulators in (c). In all panels, the system has $N_{\mathrm{sites}}=2530$ sites and $\kappa=0.01$.
@@
The corresponding bulk energy spectrum is shown in Fig. 2a as a function of $\ell$. It is symmetric with respect to zero energy as a consequence of a chiral symmetry [90] that places the model is in class $D$ of the Altland-Zirnbauer classification [92]. This feature is not essential to our findings. At the extreme values $\ell\approx 0,1$ the spectrum is gapped and we expect the bands to be topologically trivial. The reason is that at these values of $\ell$ many sites come together in real space, forming clusters of sites (see Fig. 1a and Supplemental Video). The exponential form of $f(r_{\alpha\beta})$ leads to a strong suppression of the hoppings between clusters, exponentially localizing the wave-functions to these clusters. Since exponentially localized wave-functions are topologically trivial, so the gaps are all trivial [93, 94, 95].
@@
This expectation is confirmed by computing the bulk average of the local-Chern marker [96]
$C_{\textbf{r}_{\alpha}}=2\pi\text{Im}\sum_{l}\bra{\textbf{r}_{\alpha},l}\hat{P}\left[\hat{Q}\hat{X},\hat{P}\hat{Y}\right]\hat{P}\ket{\textbf{r}_{\alpha},l},$ (2)
where $\hat{P}=\sum_{E<E_{F}}\ket{E}\bra{E}$ | is the projector onto occupied eigenstates states |$\ket{E}$ below the Fermi energy $E_{F}$, $\hat{Q}=1-\hat{P}$, and $\hat{X}$ and $\hat{Y}$ the position operators in the plane. Its average over a bulk real-space region of area $\mathcal{A}_{b}$ (see Appendix C in SM [85]), $C=(1/\mathcal{A}_{b})\sum_{\alpha\in\text{Bulk}}C_{\textbf{r}_{\alpha}}$, determines the bulk Chern number, $C$, which is approximately quantized if the occupied states are topological [70]. We find that the average of the local-Chern marker is zero when we place $E_{F}$ within any of the gaps seen at $\ell\approx 0,1$, as advertised.
@@
As we tune $\ell$, the angle- and position-dependence of the hoppings in Eq. (A Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor) allows the spectrum to evolve, see Fig. 2a. At $\ell\approx 0.28$ we observe band closings, suggesting a topological phase transition.
Topological transitions would imply the appearance of in-gap spectral weight corresponding to topological edge states in the total momentum-resolved density of states, or spectral function. We define the spectral function by projecting the real-space Hamiltonian $\mathcal{H}$ onto plane-wave states, $A(E,\mathbf{p})=\sum_{l}\bra{\textbf{p},l}\delta(\mathcal{H}-E)\ket{\textbf{p},l}$, where |$\ket{\textbf{p},l}=\sum_{\alpha}e^{i\mathbf{p}\mathbf{r}_{\alpha}}\ket{\textbf{r}_{\alpha},l}$, see Appendix B in SM [85]. This function is measurable in angle-resolved photemission spectroscopy even without translational invariance [97, 98, 99, 100, 101, 102]. Note that, unlike the crystalline case, the components of ${\bf p}$ are not restricted to the interval $\left[0,2\pi\right]$. Comparing Figs. 2b and c, at either side of the gap closing (with $\ell=0.23$ and $\ell=0.33$, respectively) we see the appearance of in-gap spectral weight, as expected for a topological transition (see also Supplemental Video).
@@
We confirm the topological nature of the in-gap states at $\ell=0.33$ by computing the local Chern marker at $E_{F}=-1.15$, see Fig. 2d. For these parameters its bulk average is $C\approx 0.98$, confirming the topological phase. These results show that we can tune the topological phase by varying $\ell$ while keeping the filling fixed. For example, at $E_{F}=-1.15$ (dashed line in Fig. 2c), a trivial aperiodic tiling leads to an aperiodic tiling with finite Chern number.
@@
We now ask, how far do the topological properties survive in parameter space, compared to the crystalline lattice model with the same Hamiltonian parameters? Because the bulk average of the local Chern marker is only approximately quantized, it is more convenient to answer this question with a local marker that is exactly an integer in the bulk. Such a marker can be obtained using the spectral localizer [103, 104, 105, 106]
=(^$\mathcal{L}=(\hat{\mathcal{H}}-E_{0}\mathbb{1})\sigma_{z}+\kappa\left[(\hat{X}-x_{0}\mathbb{1})\sigma_{x}+(\hat{Y}-y_{0}\mathbb{1})\sigma_{y}\right],$ (3)
@@
where $E_{0}$ is the reference energy at which we want to probe topology and $x_{0}$ and $y_{0}$ are chosen in the bulk of the system  [^2] In practice, we choose the coordinates $x_{0}$ and $y_{0}$ by computing the geometric mean of the vertex coordinates.. The scalar $\kappa$ is fixed to compensate the differences in scales between $\hat{\mathcal{H}}$ and the position operators $\hat{X}$, and $\hat{Y}$, see Appendix D in SM [85]. The half-signature of the localizer, $\frac{1}{2}\text{Sig}(\mathcal{L})$, defined as the half-difference between the number of positive and negative eigenvalues, is an integer that equals the bulk Chern number [103, 104, 105, 106]. In turn, the minimal eigenvalue of the spectral localizer, the localizer gap, determines the stability of a given phase, since a topological phase transition can only occur if the localizer gap closes [106], see Appendix D in SM [85].
@@
As a reference, when applied on a square lattice, the model defined in Eq. (A Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor) only has topological phases when $40<M/t<-$ 4. In contrast, we see in Fig. 3a that the phase diagram of the Hat family as a function of $\ell$ and $M/t$ extends further. The topological bands survive for much larger values of $M/t$ compared to the square lattice model, delimited by the dashed lines in Fig. 3a, especially around $\ell\approx 0.5$. Consistently, we see that the localizer gap, shown in Fig. 3b, closes at the phase boundaries.
@@
Given that $\ell$ serves as a tuning parameter for inducing topological phase transitions, one might ask whether it can also be used to control a disordered induced Anderson transition. To address this question, we introduce a random onsite potential
@@
where $V_{\alpha,ll^{\prime}}^{\phantom{}}=V_{\alpha}\delta_{ll^{\prime}}$ represent a site-dependent potential drawn from a uniform distribution in the range $[-W/2,W/2]$, where $W$ controls the disorder strength. We observe in Fig. 3b that the system remains topological for values of $W/t\leq 1$ before undergoing an Anderson transition, signaled by the closing of the average localizer gap, shown in Fig. 3d. As in crystals, for specific values of $\ell$, such as $\ell=0.8$, we observe that the system, initially trivial in the clean limit, transitions into a topological phase as $W$ increases, realizing a topological Anderson insulator  [108, 109, 110]. Interestingly, $\ell$ can be used as an additional knob to control a topological Anderson transition at fixed disorder strength, a mechanism specific to this tiling family.
Figure 4: (a) Energy spectrum as a function of $\ell$, showing a zoomed-in version of Fig. 1a including bulk and edge states. $3/4$ filling is shown as a yellow dashed line in (a). (b) Bulk trace of the local quantum metric $G$ and local Chern marker $C$ for this filling as a function of $\ell$. At small $\ell$ the sites cluster and localize the wave-functions for this filling, leading to a small $G$ with $C\approx 0$. At intermediate $\ell\approx 0.5$, $G\geq C\approx 1$, satisfying the trace inequality. For large $\ell\gtrsim 0.7$ this filling sits in the middle of a band of delocalized states with increased $G$. In both panels, the system has $N_{\mathrm{sites}}=2530$ and $M/t=-2.7$.
Geometric control of the quantum metric—The local Chern marker is the imaginary part of a more general tensor called the quantum geometric tensor [26, 27, 28, 29, 30, 31, 32] , which can also be controlled by tuning $\ell$. Its real part is known as the quantum metric [31, 29, 32], and is proportional to the real-space spread of occupied states. By expressing it as a local real-space quantity (see Appendix C in SM [85] and Refs. [71, 72])
@@
where $\hat{r}_{x}=\hat{X}$ and $\hat{r}_{y}=\hat{Y}$, we can use it to observe edge-state signatures, as shown in Fig. 2e for the topological state at $\ell=0.33$ at $E_{F}=-1.15$. We see $g_{\bf{r}_{\alpha}}$ reaches its maximum values at the edges of the system, as expected for delocalized topological edge states.
As with the bulk local Chern marker, the bulk quantum metric, is also tunable as a function of the parameter $\ell$ at a given filling. We define the bulk quantum metric by averaging Eq. (5) over a real-space region of area $\mathcal{A}_{b}$ as $G=(1/\mathcal{A}_{b})\sum_{\alpha\in\text{Bulk}}g_{{\bf r}_{\alpha}}$, see Appendix C in SM [85]. Tuning $G$ is desirable in several contexts. For instance, $G$ is known to be bounded from below by the Chern number, $G\geq C$, an inequality known as the trace inequality [111, 112, 113, 114, 115, 116]. When a band saturates this inequality it is termed vortexable or ideal [115, 116], as it is a favorable starting point to realize fractional Chern insulator states [117, 118, 119]. The quantum metric enters the superfluid stiffness [73, 74, 75, 25, 76, 77] and optical sum rules bounding topological gaps [111, 78, 79, 80]. The quantum metric depends on the real-space position, or embedding, of the orbitals [120, 121, 122, 123, 124, 125, 31, 126], controlled by $\ell$ in our system. In crystals, the embedding-dependence can be exploited to optimize responses such as the photovoltaic efficiency [124].
Fig. 4 shows the tunability of the quantum geometric tensor by plotting the trace of the local Chern marker $C$ and the local quantum metric $G$ as a function of $\ell$. For 3/4 filling (dashed yellow line in Fig. 4a) $G$ decreases as $\ell$ decreases with $C\approx 0$, see Fig. 4b. This is consistent with the fact that for these parameters the bands are narrow, topologically trivial, and states are localized due to the clustering of sites (see Supplemental Video). In contrast, at intermediate values of $\ell\approx 0.5$ the local Chern marker is approximately quantized to $C\approx 1$. The quantum metric decreases but does not saturate the trace inequality $G\geq C$, as seen in other Chern insulator models [127, 128]. At large $\ell\gtrsim 0.7$, $C$ decreases to zero while $G$ increases. This indicates that the states at this filling are more delocalized, consistent with the fact that $E_{F}$ falls within a trivial band with a relatively large spread in energy, see Fig. 4a.

### v4

# Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor

@@ Abstract
The strict geometric rules that define aperiodic tilings lead to the unique spectral and transport properties of quasicrystals, but also limit our ability to design them. In this Letter, we explore a novel example of a continuously tunable family of two-dimensional aperiodic tilings in which the underlying real-space geometry becomes a control knob of the wavefunction’s quantum geometric tensor. The real-space geometry can be used to tune into topological phases occupying an expanded phase space compared to crystals, or into a disorder-driven topological Anderson insulator. The quantum metric can also be tuned continuously, opening new routes towards tunable single- and many-body physics in aperiodic solid-state and synthetic systems.
Introduction—The electronic properties of quasicrystals can be strikingly different to those of crystals [1, 2, 3]. For example, quasicrystals often exhibit rotational symmetries forbidden in crystals and wavefunctions which are neither localized nor fully extended [1, 2, 3, 4, 5]. However, the strict mathematical rules that determine the quasicrystalline atomic arrangements and the special wavefunctions they host also restrict the possibility of tuning them. Different realizations of two-dimensional quasicrystals, such as the Penrose [6] and Ammann–Beenker [7] tilings, cannot be smoothly deformed into one another because of their different symmetries. Nonetheless, smoothly connected families of quasiperiodic structures can be generated via deformed model sets [8, 9]. Continuously tuning the electronic properties of quasicrystals is also possible by changing model parameters, as in crystals [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
Figure 1: Tiles and tilings for different values of the tile side length, defined by two side lengths, $\ell$ and 11-$\ell$. (a) shows a selection of tiles, including the Chevron, Hat, Turtle, and Comet. Their corresponding aperiodic tilings are shown in (b-e). For each $\ell$ we place two orbitals $l$ at every vertex, with onsite terms $\epsilon_{ll^{\prime}}$. Pairs of sites $\{\alpha,\beta\}$ are connected by the space and orbital dependent hopping $T_{ll^{\prime}}(\mathbf{r}_{\alpha\beta})$, see Eq. (Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor).
@@
Here we show that a newly discovered family of aperiodic tilings [70, 71] offers a real-space geometric pathway to tune the quantum geometric tensor. This family of aperiodic tilings is parametrically connected by a single real parameter $\ell$ offering a new method to geometrically tune electronic properties. This parameter determines the position of all atomic sites while preserving the quasicrystalline long-range order. Allowing the hoppings to depend on atomic positions, we show that changes in $\ell$ translate into changes of the quantum geometric tensor [72, 33, 36], expressed in terms of a real-space Chern marker [73] and quantum metric [74, 75]. Changes in the real-space Chern marker induced by $\ell$ induce phase transitions in and out of a topological phase, which exists in an expanded parameter space compared to the same model in crystals. The parameter $\ell$ also tunes the trace of the quantum metric [30, 31, 32, 33, 34], which contributes to the superfluid stiffness and kinetic inductance [76, 77, 78, 29, 79, 80], enters optical sum rules [81, 82, 83], and determines localization [84] and entanglement properties [85]. Hence, changing the real-space geometry using $\ell$ allows one to tune the quantum geometric tensor.
@@
Figure 2: Spectral, topological, and geometric properties of the model for different values of $\ell$. (a) Energy spectrum of the bulk Hamiltonian as a function of $\ell$. We obtain the bulk spectrum by excluding states whose weight inside the yellow contour in (d) and (e) is smaller than their weight outside it. Magenta dashed lines indicate the $\ell$ values used in panels (b) and (c). The yellow dashed line indicates the Fermi energy $E_{F}=-1.15t$. (b) Spectral function $\mathcal{A}(E,{\bf p})$ for $\ell=0.23$, showing a full gap near 1/4 filling, attained near the energy indicated by the white dashed line. (c) Spectral function $\mathcal{A}(E,{\bf p})$ for $\ell=1/3$, showing the presence of edge states within bulk gaps. (d) Local Chern marker $\mathcal{C}_{{\bf r}_{i}}$ Eq. (2), evaluated at the energy indicated by the white dashed line in (c) for $\ell=1/3$. The yellow solid line outlines the bulk region used to perform the averaging. In this region, the average value of the local Chern marker is $0.98$. (e) Local quantum metric $g_{\bf r}$, Eq. (5), computed at the $E_{F}$ indicated by the dashed line in (c) for $\ell=1/3$. We observe that the local quantum metric takes large values at the edges, indicating delocalized edge states. In all the panels the system has 25302530 sites and M/t=2.7M/t=-2.7. In panels (b) and (c) we used the kernel polynomial method [86] to compute the spectral function with 512 moments.
@@
Geometric control of topology—Fig. 1a shows the Hat family of aperiodic tilings, parametrized by a real number $\ell\in\left[0,1\right]$. It parametrizes the two independent lengths of parallel sides, $a=\ell$ and $b=1$-$\ell$. By tuning $\ell$, one can continuously interpolate between different tiles in the family. The Hat tiling $\ell=1/(1+\sqrt{3})$] made headlines as the first aperiodic monotile [70, 71]. Other named tiles are the Specter $\ell=1/2$], the Turtle $\ell=\sqrt{3}/(1+\sqrt{3})$], the Chevron $\ell=0$], and the Comet $\ell=1$] which sets our unit of length. A previous Letter has shown that a non-bipartite version of the Hat tiling shares similar electron properties with those of graphene [25] when all the hoppings are set to be equal, inherited from its sixfold rotational symmetry [87]. Here, we focus on the bipartite version of these tilings, and describe their construction in the Supplemental Material (SM) [88], by adapting the code published along with Ref. [89].
@@
Figure 3: Topological phase diagram. (a) Bulk half-signature of the spectral localizer at $E_{0}=E_{F}$ as a function of both the geometrical parameter $\ell$ and the value of the onsite term $M/t$. $E_{F}$ is fixed at $3/4$ filling out of a total of $N_{\mathrm{sites}}\times N_{\mathrm{orbs}}=2530\times 2$ states. We observe topological phases for a wider parameter region compared to the square-lattice crystalline model, whose phase boundaries are delimited by the black dashed lines. (b) Localizer gap corresponding to (a). The gap in the spectrum of the spectral localizer closes at the boundary between the topological and trivial phases of panel (a). (c) Bulk half-signature of the spectral localizer averaged over 100 disorder realizations at $E_{0}=E_{F}$ and $M/t=0.5$, as a function of $\ell$ and onsite disorder strength $W/t$. We observe two types of behavior as we increase $W/t$ for fixed $\ell$: either the clean system is topological and transitions into a trivial Anderson insulator (e.g. at $\ell=0.5$), or a trivial clean system transitions into a topological phase, to later become a trivial Anderson insulator (e.g. at $\ell=0.8$ or $\ell=0.2$). (d) The localizer gap vanishes between the topological and trivial Anderson insulators in (c). In all panels, the system has $N_{\mathrm{sites}}=2530$ sites and $\kappa=0.01$. Panels (b) and (d) are plotted using a logarithmic colormap.
@@
The corresponding bulk energy spectrum is shown in Fig. 2a as a function of $\ell$. It is symmetric with respect to zero energy as a consequence of a chiral symmetry [93] that places the model in class $D$ of the Altland-Zirnbauer classification [95]. This feature is not essential to our findings. At the extreme values $\ell\approx 0,1$ the spectrum is gapped and we expect the bands to be topologically trivial. The reason is that at these values of $\ell$ many sites come together in real space, forming clusters of sites (see Fig. 1a and Supplemental Video). The exponential form of $f(r_{\alpha\beta})$ leads to a strong suppression of the hoppings between clusters, exponentially localizing the wavefunctions to these clusters. Since exponentially localized wavefunctions are topologically trivial, so are the gaps [96, 97, 98].
@@
This expectation is confirmed by computing the bulk average of the local Chern marker [99]
[…]
@@
As we tune $\ell$, the angle- and position-dependence of the hoppings in Eq. (Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor) allows the spectrum to evolve, see Fig. 2a. At $\ell\approx 0.28$ we observe gap closings across the whole energy range, suggesting potential topological phase transitions.
[…]
@@
We confirm the topological nature of the in-gap states at $\ell=0.33$ by computing the local Chern marker at $E_{F}=-1.15t$, corresponding to 1/4 filling, see Fig. 2d. For these parameters its bulk average is $C\approx 0.98$, confirming the topological phase. These results show that we can tune the topological phase by varying $\ell$ while keeping the filling fixed. For example, at $E_{F}=-1.15t$ (yellow dashed line in Fig. 2a), by varying $\ell$, a state that is initially trivial becomes topological.
@@
We now ask, how far do the topological properties survive in parameter space, compared to the crystalline lattice model with the same Hamiltonian parameters? To answer this question it is convenient to use the spectral localizer index, which is obtained from the spectral localizer [106, 107, 108, 109]
$\mathcal{L}=(\hat{\mathcal{H}}-E_{0}\mathbb{1})\sigma_{z}+\kappa\left[(\hat{X}-x_{0}\mathbb{1})\sigma_{x}+(\hat{Y}-y_{0}\mathbb{1})\sigma_{y}\right],$ (3)
@@
where $E_{0}$ is the reference energy at which we want to probe topology and $x_{0}$ and $y_{0}$ are chosen in the bulk of the system  [^1] In practice, we choose the coordinates $x_{0}$ and $y_{0}$ by computing the geometric mean of the vertex coordinates.. The scalar $\kappa$ is fixed to compensate the differences in scales between $\hat{\mathcal{H}}$ and the position operators $\hat{X}$, and $\hat{Y}$, see SM [88]. The half-signature of the localizer, $\frac{1}{2}\text{Sig}(\mathcal{L})$, defined as the half-difference between the number of positive and negative eigenvalues, is an integer that equals the bulk Chern number [106, 107, 108, 109].
Although the local Chern marker and the spectral localizer provide equivalent topological information [111], the spectral localizer additionally quantifies the robustness of the topological phase. In particular, its minimal eigenvalue, known as the localizer gap, serves as a measure of phase stability, since a topological phase transition can only occur when this gap closes [109, 88]. Because the localizer index does not require calculating projectors, topological phase diagrams can be computed numerically more efficiently than the local Chern marker average [109].
@@
As a reference, when applied on a square lattice, the model defined in Eq. (Family of Aperiodic Tilings with Tunable Quantum Geometric Tensor) only has topological phases when 4<M/t<0-4<M/t<0. In contrast, we see in Fig. 3a that the phase diagram of the Hat family as a function of $\ell$ and $M/t$ extends further. The topological bands survive for much larger values of $M/t$ compared to the square lattice model, delimited by the dashed lines in Fig. 3a, especially around $\ell\approx 0.5$  [^2] The edge statistics explain why topological regions widen around $\ell=0.5$, and why the phase diagram is asymmetric about this point. For generic $\ell$ there are $N_{s}$ short and $N_{l}$ long bonds, satisfying $N_{s}>N_{l}$ for $\ell<0.5$, $N_{s}<N_{l}$ for $\ell>0.5$, and $N_{\rm s}=N_{l}$ for $\ell=0.5$. Since the average spacing $\bar{a}$ in $f(r_{\alpha\beta})$ increases with $\ell$, this imbalance yields an asymmetric hopping distribution. At $\ell=0.5$ all hoppings are equal, minimizing detrimental bond disorder [179, 180] and leaving only connectivity disorder, as in Weaire–Thorpe models [181]. Away from $\ell=0.5$, the two competing hopping strengths associated to $N_{s}$ and $N_{l}$ drive the system toward trivial, Wannierizable dimerized limits at $\ell\to 0,1$, leaving $\ell=0.5$ as the optimal value.. Consistently, we see that the localizer gap, shown in Fig. 3b, closes at the phase boundaries.
@@
Given that $\ell$ serves as a tuning parameter for inducing topological phase transitions, one might ask whether it can also be used to control a disorder-induced topological phase transition. To address this question, we introduce a random onsite potential
@@
where $V_{\alpha,ll^{\prime}}^{\phantom{}}=V_{\alpha}\delta_{ll^{\prime}}$ represents a site-dependent potential drawn from a uniform distribution in the range [W/2,W/2][-W/2,W/2], where $W$ controls the disorder strength. We observe in Fig. 3c that the system remains topological for disorder strengths up to $W/t\leq 1$, before undergoing an Anderson transition, signaled by the closing of the average localizer gap shown in Fig. 3d. As in crystals, for certain values of $\ell$ (e.g., $\ell=0.8$), the system, initially trivial in the clean limit, enters a topological phase as $W$ increases  [^3] This phase is sometimes referred to as topological Anderson insulator phase even though it is connected to the disorder-free limit, indicating that disorder can enlarge the topological region in parameter space [114, 115, 116]. Interestingly, $\ell$ serves as an additional tuning knob that enables a topological phase transition at fixed disorder strength, a mechanism specific to this tiling family.
Figure 4: (a) Energy spectrum as a function of $\ell$, showing a zoomed-in version of Fig. 2a including bulk and edge states. $3/4$ filling is shown as a yellow dashed line in (a). (b) Bulk trace of the local quantum metric $G$ and local Chern marker $C$ for this filling as a function of $\ell$. At small $\ell$ the sites cluster and localize the wavefunctions for this filling, leading to a small $G$ with $C\approx 0$. At intermediate $\ell\approx 0.5$, $G\geq C\approx 1$, satisfying the trace inequality. For large $\ell\gtrsim 0.7$ this filling sits in the middle of a band of delocalized states with increased $G$. In both panels, the system has $N_{\mathrm{sites}}=2530$ and M/t=2.7M/t=-2.7.
Geometric control of the quantum metric—The local Chern marker is the imaginary part of a more general tensor called the quantum geometric tensor [30, 31, 32, 33, 34, 35, 36] , which can also be controlled by tuning $\ell$. Its real part is known as the quantum metric [35, 33, 36], and is proportional to the real-space spread of occupied states. By expressing it as a local real-space quantity (see SM [88] and Refs. [74, 75])
@@
where $\hat{r}_{x}=\hat{X}$ and $\hat{r}_{y}=\hat{Y}$, we can use it to observe edge-state signatures, as shown in Fig. 2e for the topological state at $\ell=0.33$ at $E_{F}=-1.15t$, approximately at 1/4 filling. We see $g_{\bf{r}_{\alpha}}$ reaches its maximum values at the edges of the system, as expected for delocalized topological edge states. In contrast, for the trivial phase $g_{\bf{r}_{\alpha}}$ is peaked around clusters of sites, consistent with a trivial phase, see SM [88].
As with the bulk local Chern marker, the bulk quantum metric, is also tunable as a function of the parameter $\ell$ at a given filling. We define the bulk quantum metric by averaging Eq. (5) over a real-space region of area $\mathcal{A}_{b}$ as $G=(1/\mathcal{A}_{b})\sum_{\alpha\in\text{Bulk}}g_{{\bf r}_{\alpha}}$, see SM [88]. Tuning $G$ is desirable in several contexts. For instance, $G$ is known to be bounded from below by the Chern number, $G\geq C$, an inequality known as the trace inequality [117, 118, 119, 120, 121, 122]. When a band saturates this inequality it is termed vortexable or ideal [121, 122], as it is a favorable starting point to realize fractional Chern insulator states [123, 124, 125]. The quantum metric enters the superfluid stiffness [76, 77, 78, 29, 79, 80] and optical sum rules bounding topological gaps [117, 81, 82, 83]. The quantum metric depends on the real-space position, or embedding, of the orbitals [126, 127, 128, 129, 130, 131, 35, 132], controlled by $\ell$ in our system. In crystals, the embedding-dependence can be exploited to optimize responses such as the photovoltaic efficiency [130].
Fig. 4 shows the tunability of the quantum geometric tensor by plotting the trace of the local Chern marker $C$ and the local quantum metric $G$ as a function of $\ell$. For 3/4 filling (dashed yellow line in Fig. 4a) $G$ decreases as $\ell$ decreases with $C\approx 0$, see Fig. 4b. This is consistent with the fact that for these parameters the bands are narrow, topologically trivial, and states are localized due to the clustering of sites (see Supplemental Video). In contrast, at intermediate values of $\ell\approx 0.5$ the local Chern marker is approximately quantized to $C\approx 1$. The quantum metric decreases but does not saturate the trace inequality $G\geq C$, as seen in other Chern insulator models [133, 134]. At large $\ell\gtrsim 0.7$, $C$ decreases to zero while $G$ increases. This indicates that the states at this filling are more delocalized, consistent with the fact that $E_{F}$ falls within a trivial band with a relatively large spread in energy, see Fig. 4a. Saturating the trace inequality in the topological region may be possible by optimizing within the large parameter space of fillings and model parameters, an interesting direction for future work.